# algorithmic

### v1

@@ A.1.2 Fixed Budget
From this, successive elimination algorithms emerged Audibert et al. (2010). In these algorithms, the agent pulls every arm at each round then sees if any arms can be eliminated with high probability. The agent then stops when only one arm is remaining. This approach was proven optimal in Carpentier & Locatelli (2016) for compactly supported distributions and provides inspiration for our low-complexity algorithm.
Cost has been introduced into the fixed budget case as a natural extension of the budget. Specifically, non-deterministic cost has been explored in the fixed budget setting for both discrete Ding et al. (2013) and continuous cost Xia et al. (2016). In both cases, the density functions of the costs are compactly supported. Fixed budget BAI can be thought of as the dual problem to our problem setting. Instead of trying to maximize the confidence subject to a hard cost constraint, we try to minimize the cost subject to a hard confidence constraint.
@@ A.2 Regret Minimization
Regret minimization is concerned with minimizing the number of pulls of suboptimal arms or equivalently maximizing the number of times the arm with the highest reward is pulled. Confidence interval based algorithms have proven effective for this setting Lai & Robbins (1985); Lai (1987). This led to fruitful work in finding tighter confidence intervals utilizing Kullback-Liebler divergence Cappé et al. (2013).
@@ A.2.1 Explore Then Commit
Related to BAI, there is a category of regret minimization algorithms that follow an explore-then-commit approach Garivier et al. (2016). In these algorithms, the agent first attempts to identify the best arm as in BAI, then the agent commits to only pulling the identified best arm. However, it was shown in Degenne et al. (2019b) that any ETC algorithm that first utilizes an optimal BAI algorithm will not achieve the optimal regret bounds from Lai & Robbins (1985); Burnetas & Katehakis (1996).
@@ A.2.2 Regret Minimization With Cost
Appendix B Exponential Family
Each of the Bernoulli, Poisson, and Gaussian distributions can be realized as part of the natural exponential family by considering the following choice of parameters:
Appendix C Proofs for the Lower Bound Theorem and Related Corollaries
C.1 Proof of the General Lower Bound Theorem. 1
@@ Appendix C Proofs for the Lower Bound Theorem and Related Corollaries
Similar to the proof of Theorem 1 in Garivier & Kaufmann (2016), we first invoke the ’transportation’ lemma (Kaufmann et al., 2016a, Lemma 1), (Garivier & Kaufmann, 2016, Theorem 1) to build a relationship between the expected number of draws $N_{a}(t)$ and the Kullback-Leibler divergence from bandit instance $\boldsymbol{\mu}$ to another bandit model $\boldsymbol{\lambda}$ which has a different best arm: for any $\boldsymbol{\lambda}\in\mathcal{S}$ such that $a^{*}(\boldsymbol{\lambda})\neq a^{*}(\boldsymbol{\mu})$, we have:
@@ C.2 Proof of Lemma 1
C.3 Jensen-Shannon Divergence Form of Lower Bound
@@
C.5 Lower Bound on Another Simplified MAB Model
@@
Appendix D Computing the Optimal Proportion $\boldsymbol{w}^{*}$
D.1 Characterization of Optimal Proportion $\boldsymbol{w}^{*}$
@@
D.2 Computing $\boldsymbol{w}^{*}$ Efficiently
@@
D.3 Proofs for Lemma. 2 and Theorem. 4
@@
Appendix E Proof Roadmap of Theorem. 2
We first show that the empirical cost proportion converges to the optimal proportion $\boldsymbol{w}^{*}$ in Section. F and in Theorem. 5. Then, we prove a weaker version of Theorem. 2 which states that the cost performance of $\mathsf{CTAS}$ matches the lower bound asymptotically with probability $1$, i.e., Theorem. 6 in Sec. G. Finally, we prove Theorem. 2 in Sec. H.
Appendix F Asymptotic Convergence of Cost Proportions $\boldsymbol{\widehat{w}}(t)$ for $\mathsf{CTAS}$
@@ F.1 Proof of Theorem 5
which implies $\lim_{t\to\infty}\widehat{w}_{a}(t)\to w^{*}_{a}$ for every $\omega\in\mathcal{E}$.
F.2 Proof of Lemma. 5
@@
Appendix G Asymptotic Cumulative Cost Upper Bound for $\mathsf{CTAS}$
@@
G.2 Asymptotic Expectation Optimality
@@
Appendix H Asymptotic Cost Optimality for Chernoff-Overlap in Two-armed Gaussian Bandits
@@
Appendix I $\delta$-PAC Analysis for $\mathsf{CTAS}$ and $\mathsf{CO}$

### v2

@@ A.1.2 Fixed Budget
From this, successive elimination algorithms emerged (Audibert & Bubeck, 2010). In these algorithms, the agent pulls every arm at each round then sees if any arms can be eliminated with high probability. The agent then stops when only one arm is remaining. This approach was proven optimal in Carpentier & Locatelli (2016) for compactly supported distributions and provides inspiration for our low-complexity algorithm.
Cost has been introduced into the fixed budget case as a natural extension of the budget. Specifically, non-deterministic cost has been explored in the fixed budget setting for both discrete (Ding et al., 2013) and continuous cost (Xia et al., 2016). In both cases, the density functions of the costs are compactly supported. Fixed budget BAI can be thought of as the dual problem to our problem setting. Instead of trying to maximize the confidence subject to a hard cost constraint, we try to minimize the cost subject to a hard confidence constraint.
@@ A.2 Regret Minimization
Regret minimization is concerned with minimizing the number of pulls of suboptimal arms or equivalently maximizing the number of times the arm with the highest reward is pulled. Confidence interval based algorithms have proven effective for this setting (Lai & Robbins, 1985; Lai, 1987). This led to fruitful work in finding tighter confidence intervals utilizing Kullback-Liebler divergence (Cappé et al., 2013).
@@ A.2.1 Explore Then Commit
Related to BAI, there is a category of regret minimization algorithms that follow an explore-then-commit approach (Garivier et al., 2019). In these algorithms, the agent first attempts to identify the best arm as in BAI, then the agent commits to only pulling the identified best arm. However, it was shown in Degenne et al. (2019b) that any ETC algorithm that first utilizes an optimal BAI algorithm will not achieve the optimal regret bounds from Lai & Robbins (1985); Burnetas & Katehakis (1996). A further refined trade-off between best arm identification and regret minimization is studied in Zhang & Ying (2023).
@@ A.2.2 Regret Minimization With Cost
Appendix B Proofs for the Lower Bound Theorem and Related Corollaries
B.1 Proof of the General Lower Bound Theorem. 1
@@ Appendix B Proofs for the Lower Bound Theorem and Related Corollaries
Similar to the proof of Theorem 1 in Garivier & Kaufmann (2016), we first invoke the ’transportation’ lemma (Kaufmann et al., 2016, Lemma 1), (Garivier & Kaufmann, 2016, Theorem 1) to build a relationship between the expected number of draws $N_{a}(t)$ and the Kullback-Leibler divergence from bandit instance $\boldsymbol{\mu}$ to another bandit model $\boldsymbol{\lambda}$ which has a different best arm: for any $\boldsymbol{\lambda}\in\mathcal{S}$ such that $a^{*}(\boldsymbol{\lambda})\neq a^{*}(\boldsymbol{\mu})$, we have:
@@ B.2 Proof of Lemma 1
B.3 Jensen-Shannon Divergence Form of Lower Bound
@@
B.5 Lower Bound on Another Simplified MAB Model
@@
Appendix C Computing the Optimal Proportion $\boldsymbol{w}^{*}$
C.1 Characterization of Optimal Proportion $\boldsymbol{w}^{*}$
@@
C.2 Computing $\boldsymbol{w}^{*}$ Efficiently
@@
C.3 Proofs for Lemma. 2 and Theorem. 4
@@
Appendix D Proof Roadmap of Theorem. 2
We first show that the empirical cost proportion converges to the optimal proportion $\boldsymbol{w}^{*}$ in Section. E and in Theorem. 5. Then, we prove a weaker version of Theorem. 2 which states that the cost performance of $\mathsf{CTAS}$ matches the lower bound asymptotically with probability $1$, i.e., Theorem. 6 in Sec. F. Finally, we prove Theorem. 2 in Sec. G.
Appendix E Asymptotic Convergence of Cost Proportions $\boldsymbol{\widehat{w}}(t)$ for $\mathsf{CTAS}$
@@ E.1 Proof of Theorem 5
which implies $\lim_{t\to\infty}\widehat{w}_{a}(t)\to w^{*}_{a}$ for every $\omega\in\mathcal{E}$. ∎
E.2 Proof of Lemma. 5
@@
Appendix F Asymptotic Cumulative Cost Upper Bound for $\mathsf{CTAS}$
@@
F.2 Asymptotic Expectation Optimality
@@
Appendix G Asymptotic Cost Optimality for Chernoff-Overlap in Two-armed Gaussian Bandits
@@
Appendix H $\delta$-PAC Analysis for $\mathsf{CTAS}$ and $\mathsf{CO}$